James Robert, a 72-year-old african american make, is admitted to university hospital with dehydration and acute respiratory failure. Mr. Roberts' past medical history includes hypertension and chronic obstructive pulmonary disease (COPD). His current medications include clonidine (anti-hypertensive), prednisone (corticosteroid), and albuterol (bronchodilator).
Since the death of his wife one year ago, Mr.Roberts lives alone and does not cook for himself. Per patient's daughter, Mr. Roberts receives home-delivered meals from the food assistance program, Meals on Wheels. In addition to the one meal per day from Meals on Wheels, he relies on ready-to-eat foods, such as cereal, toast, canned fruit, cheese, processed deli meats (e.g. bologna), and snacks, such as potato chips and cookies. Due to ill-fitting dentures, Mr.Roberts does not eat freah fruits and vegetables or unprocessed meats often.
Mr.Roberts has been placed on mechanical ventilator due to his repiratory failure. Dr.Connors, the physician, wants to start nutrition support in the next 24 hours. As his nurse, you listen for bowel sounds and thhey are positive in all four quadrants of the abdomen.
The following information was recorded on admission:
Height: 5'10" Weight: 155 lbs Usual Weight (3 months ago): 170 lbs

Q. Calculate Mr.Robert's ideal body weight. Assume Mr.Robert's has a medium frame.
Mr.Robert's ideal body weight is:
a. 150 lbs
b. 160 lbs
c. 166 lbs
d. 175 lbs
Show the formula, calculations, and include the unit of measure with the answer.
Answer #1

The formula for Ideal body weight (IBW):

  • IBW = 50 kgs + 2.3kgs for each inch above 5 feet (for male) ----- (1)
  • IBW = 45.5kgs + 2.3kgs for each inch above 5 feet (for female)

Calculation:

  1. Height of the patient - 5 feet 10 inch
  2. Inserting values in formula number (1)
  3. IBW = 50 kgs + 2.3 kgs ( 10 inch )
  4. IBW = 50 + 23
  5. IBW = 73 kgs
  6. Converting kilos into pounds (lbs) - the conversion factor is 1 kg = 2.2 lbs
  7. Therefore IBW = 73 * 2.2 = 160.6 lbs ( option b)